Can you please help me solve for x

500/(x+5)+5=500/x

Answer by checkley71(8403)

500/(X+5)+5=500/X
500/(X+5)-500/X=-5
[500*X-500(X+5)]/X(X+5)=-5
(500X-500X-2500)/(X^2+5X)=-5 CROSS MULTIPLY
-5(X^2+5X)=-2500
-5X^2-25X+2500=0
-5(X^2+5X-500)=0
-5(X+25)(X-20)=0
X+25=0
X=-25 ANSWER.
X-20=0
X=20 ANSWER.